An orderly is pushing a gurney with a patient on it. The mass of the orderly is 85 kg, the mass of the gurney is 20 kg, and the mass of the patient is 50 kg. The orderly exerts a backward force of 100 N on the floor. What acceleration is produced, assuming the friction in the wheels is negligible? 
Answer #1

total force on the system = 100 N

total mass of the system = 85 + 20 + 50

= 155 kg

acceleration = F/m

= 100 / 155 = 0.645 m/s^2

so the acceleration produced = 0.645 m/s^2